score: U ot1 p Bus Econ 3.3.29 in order to accumulate enou Type the amount in the account s ty money for a down payment on a house, .co-deposts S20,pe morth me an account paye 96%con nded morni,paymers . * heend Round to the nearest dollar)
Answer #1

Future value

Using future value function in M S excel

fv(rate,nper,pmt,pv,type) rate = 6/12 = .5% nper = 12*4 = 48 pmt = 209 pv =0 type =0

FV(0.5%,48,209,0,0)

($11,306)